Show that the series Σ x/(1+n^2.x^2) is uniformly convergent using Weierstrass m test


1
Expert's answer
2021-07-16T13:41:25-0400

Since for each "n\in N" and "x\in \left[ 1,\alpha \right]" , we have "1+{ n }^{ 2 }{ x }^{ 2 }\ge { n }^{ 2 }\ ,\ x\le \alpha" ,then


"0\le \frac { x }{ 1+{ n }^{ 2 }{ x }^{ 2 } } \le \frac { \alpha }{ { n }^{ 2 } }"

 .

Setting "{ f }_{ n }(x)=\frac { x }{ 1+{ n }^{ 2 }{ x }^{ 2 } } ,\ { M }_{ n }=\frac { \alpha }{ { n }^{ 2 } }" we have for each "n\in N" and "x\in \left[ 1,\alpha \right]" : 


"\left| { f }_{ n }(x) \right| ={ f }_{ n }(x)\le { M }_{ n }" . The series "\sum _{ n=1 }^{ \infty }{ { M }_{ n }\ } =\alpha \sum _{ n=1 }^{ \infty }{ \frac { 1 }{ { n }^{ 2 } } }" converges, because the series 



"\sum _{ n=1 }^{ \infty }{ \frac { 1 }{ { n }^{ 2 } } }" is p-series (p=2). By the Weierstrass M-Test , the series


"\sum _{ n=1 }^{ \infty }{ { f }_{ n }(x)= } \sum _{ n=1 }^{ \infty }{ \frac { x }{ 1+{ n }^{ 2 }{ x }^{ 2 } } }"


converges uniformly in "\left[ 1,\alpha \right]" .